# references.bib

### v3

@@ IV-C Evaluation on Compression Performance
DLRM End-to-end Performance Speed-up Figure 1 shows the breakdown of DLRM training on our cluster with 32 A100 GPUs. Our compression accelerates all-to-all communication in forward propagation which takes 31.3% proportion of the whole training time. On Criteo Kaggle dataset, our proposed compressor achieves 6.22$\times$ overall speed-up in communication and 1.30$\times$ in end-to-end training, by reducing all-to-all in forward propagation cost to 5.03%. On Criteo Terabytes dataset, our compressor achieves 8.6$\times$ and 1.38$\times$ in all-to-all communication and end-to-end training, respectively.
@@
Compression Performance across Training Phases. Based on Figures 10(b) and 11(b), we observe that our compressor performs effectively throughout all training phases, maintaining a consistently high compression ratio. Implementing an error bound decay to enhance the quality of compressed data results in only a slight decrease in the compression ratio. This stable compression ratio can be attributed to two main factors. Firstly, the Vector-based LZ encoder, which relies on pattern matching, maintains its effectiveness due to the presence of identifiable patterns in each batch, a factor that remains constant regardless of the training phase. Secondly, the uniformity in data distribution across the training, as depicted in Figure 14, ensures consistent compression performance.
@@ IV-D Evaluation of Compression Optimization
LZ Fine-tuning. First, we evaluate our optimized compressor with different LZ fine-tuning window sizes. Table VI shows the compression ratio and throughput on DLRM data with varying window sizes. Generally, larger window sizes result in more pattern matches. For the Criteo Terabyte dataset, we observe that the overall compression ratio with the window sizes of 128 and 255 are 3.9$\times$ and 5.2$\times$ higher, respectively, compared to the baseline window size of 32. On the Criteo Kaggle dataset, the difference between the window sizes of 128 (1.52$\times$) and 255 (1.54$\times$) is negligible. In instances of small batch sizes, our vector-based LZ is less beneficial with the increase in window size. This is due to the proportion between the sliding window size and the volume of data. In the model of Criteo Kaggle dataset, for example, $BatchSize$ is 128 by default and can be fully covered by one sliding window. The scaled compression ratios of the Criteo Terabyte dataset demonstrate that increasing the window size does not linearly increase the compression ratio. This phenomenon is attributed to the unbalanced frequency of queries. EMB vectors with very high frequency can be matched with an appropriate window size, whereas EMB vectors with low frequency require a window size that exceeds hardware limitations and is inefficient.
@@
Buffer Optimization. Second, we evaluate our buffer optimization in both compression and decompression processes. We split EMB vectors into chunks corresponding to the $RANK$ in distributed training, indicative of training scalability. We use the chunk size ranging from 2 to 16. Figure 15 illustrates compression speedup with various data sizes. Specifically, our optimization achieves a maximum speedup of 2.04$\times$ in scenarios involving large scale, with performance enhancement evident as $RANK$ increases.
@@
Figure 15: Speedup of our buffer optimization.
@@ V Related Work
Low-precision Training. Exploring low-precision data types, as proposed by Rouhani et al. with the new datatype MX[0.02threshold] and mixed-precision strategies by Yang et al.[yang2020mixedprecision], is another direction of research. These methods, though direct, offer limited compression benefits due to their fixed compression ratio and lack fine-tuning capabilities, providing a coarse granularity control over data compression. Our strategy, in contrast, allows for a smooth adjustment of error bounds, offering a more flexible and efficient solution to communication reduction in DLRM training.
@@ VI Conclusion and Future Work
In this paper, we introduce a method that employs error-bounded lossy compression to reduce the communication data size and accelerate DLRM training. We develop a novel error-bounded lossy compression algorithm to achieve hybrid by hybridizing our optimized LZ encoder and entropy encoder on GPU. Moreover, we introduce a dual-level adaptive strategy for error-bound adjustment, spanning both table-wise and iteration-wise aspects, to balance the compression benefits with the potential impacts on accuracy. We further optimize our compressor for PyTorch tensors on GPUs, minimizing compression overhead. Evaluation shows that our method achieves a 8.6$\times$ all-to-all communication speedup and 1.38$\times$ end-to-end training speedup with a minimal accuracy impact.
In the future, we plan to further refine our system to reduce compression overhead by employing strategies like kernel fusion on GPUs and seamlessly integrating (de)compression processes with communication libraries such as NCCL. Additionally, we aim to develop a more advanced and automated approach for offline selection of a fixed global error bound and for online error-bound adjustments.
@@ Acknowledgment
The work is supported by the Meta Research Award for “AI System Hardware/Software Codesign.” The work is supported by the National Science Foundation (Grant Nos. 2312673, 2311876, 2247080, 2303064, 2326494, and 2326495).

### v4

@@ IV-C Evaluation on Compression Performance
DLRM End-to-end Performance Speed-up Figure 12 shows the breakdown of DLRM training with lossy compression on our cluster with 32 A100 GPUs. Our compression accelerates all-to-all communication in forward propagation which takes 31.3% proportion of the whole training time. On the Criteo Kaggle dataset, our proposed compressor achieves 6.22$\times$ overall speed-up in communication and 1.30$\times$ in end-to-end training, by reducing all-to-all in forward propagation cost to 5.03%. On the Criteo Terabytes dataset, our compressor achieves 8.6$\times$ and 1.38$\times$ in all-to-all communication and end-to-end training, respectively.
@@
Compression Performance across Training Phases. Based on Figures 9(b) and 10(b), we observe that our compressor performs effectively throughout all training phases, maintaining a consistently high compression ratio. Implementing an error-bound decay to enhance the quality of compressed data results in only a slight decrease in the compression ratio. This stable compression ratio can be attributed to two main factors. Firstly, the Vector-based LZ encoder, which relies on pattern matching, maintains its effectiveness due to the presence of identifiable patterns in each batch, a factor that remains constant regardless of the training phase. Secondly, the uniformity in data distribution across the training, as depicted in Figure 14, ensures consistent compression performance.
@@ IV-D Evaluation of Compression Optimization
LZ Fine-tuning. First, we evaluate our optimized compressor with different LZ fine-tuning window sizes. Table VI shows the compression ratio and throughput on DLRM data with varying window sizes. Generally, larger window sizes result in more pattern matches. For the Criteo Terabyte dataset, we observe that the overall compression ratio with the window sizes of 128 and 255 are 3.9$\times$ and 5.2$\times$ higher, respectively, compared to the baseline window size of 32. On the Criteo Kaggle dataset, the difference between the window sizes of 128 (1.52$\times$) and 255 (1.54$\times$) is negligible. In instances of small batch sizes, our vector-based LZ is less beneficial with the increase in window size. This is due to the proportion between the sliding window size and the volume of data. In the model of the Criteo Kaggle dataset, for example, $BatchSize$ is 128 by default and can be fully covered by one sliding window. The scaled compression ratios of the Criteo Terabyte dataset demonstrate that increasing the window size does not linearly increase the compression ratio. This phenomenon is attributed to the unbalanced frequency of queries. EMB vectors with very high frequency can be matched with an appropriate window size, whereas EMB vectors with low frequency require a window size that exceeds hardware limitations and is inefficient.
@@
Buffer Optimization. Second, we evaluate our buffer optimization in both compression and decompression processes. We split the EMB vectors into chunks, with the number of chunks equal to the RANK in distributed training, which reflects the scalability of the training process. Figure 15 illustrates the compression speedup across different EMB vector sizes, with the number of chunks ranging from 2 to 16. The term ’single_comp’ denotes our solution. The results indicate that our design achieves higher speedups with an increased number of chunks. According to our evaluation, our optimizations achieve a maximum speedup of 2.04$\times$.
@@
Figure 15: Normalized time of our work with and without buffer optimization for different EMB vector sizes. ‘single_comp‘ denotes our solution, while the chunk number indicates how many chunks the original EMB vector is equally partitioned and compressed.
@@ V Related Work
Low-precision Training. Exploring low-precision data types, as proposed by Rouhani et al. with the new datatype MX[0.02threshold] and mixed-precision strategies by Yang et al.[yang2020mixedprecision], is another direction of research. These methods, though direct, offer limited compression benefits due to their fixed compression ratio and lack of fine-tuning capabilities, providing a coarse granularity control over data compression. Our strategy, in contrast, allows for a smooth adjustment of error bounds, offering a more flexible and efficient solution to communication reduction in DLRM training.
@@ VI Conclusion and Future Work
In this paper, we introduce a method that employs error-bounded lossy compression to reduce the communication data size and accelerate DLRM training. We develop a novel error-bounded lossy compression algorithm to achieve hybrid by hybridizing our optimized LZ encoder and entropy encoder on GPU. Moreover, we introduce a dual-level adaptive strategy for error-bound adjustment, spanning both table-wise and iteration-wise aspects, to balance the compression benefits with the potential impacts on accuracy. We further optimize our compressor for PyTorch tensors on GPUs, minimizing compression overhead. The evaluation shows that our method achieves an 8.6$\times$ all-to-all communication speedup and 1.38$\times$ end-to-end training speedup with a minimal accuracy impact.
In the future, we plan to further refine our system to reduce compression overhead by employing strategies like kernel fusion on GPUs and seamlessly integrating (de)compression processes with communication libraries such as NCCL. Additionally, we aim to develop a more advanced and automated approach for offline selection of a fixed global error-bound and for online error-bound adjustments.
@@ Acknowledgment
The work is supported by the Meta Research Award for “AI System Hardware/Software Codesign.” The work is supported by the National Science Foundation (Grant Nos. 2312673, 2311876, 2247080, 2303064, 2326494, and 2326495). The work is also supported by the National Natural Science Foundation of China (Grant Nos. 62032023 and T2125013) and the Innovation Funding of ICT, CAS (Grant No. E461050).